AGw Commissioning with IRTC - 12-13 November 2008

Observers: J. Hill (first half), D. Miller and A. Rakich in Tucson plus J. Storm (second half) in Potsdam
Software Support: T. Edgin (first half)
Telescope Operator: W. Wack

Summary

Had some minor problems with the new TCS build. Then took tracking test data and checked the Z22 coefficient in fairly poor seeing. Jesper took some more AGW calibration data at the end of the night.

Details (Times reported in MST below)

New Stuff

JMH updated the global offsets in the M2 collimation model this afternoon, so it now INCLUDES the -11 arcsec of M2 tip that you applied last night. Specifically: -11 arcsec of Tip changed Y by +0.562 mm and changed RX by -108.02 arcsec.

I did not adjust the collimation models to include the coordinated Y shift, although I think that I've understood how to do that in the temperature terms if we need to.

A new build of TCS is going up this afternoon.

Dave A. reports that they have identified the problem with the rotator dropping to IDLE, and they have increased the torque limit to prevent this problem from happening when it gets cold.

Note that the new TCS build over-wrote the 09-Nov-2008 version of the Zernike coefficients in AIP_L_RM_sub11.cfg, so we were using the older value in GCS (but not IDL) for this night.

Open

Time in MST

19:30 After a 1/2 hour delay as people finished work on HBS, and then the letting the DIMM guy restart his system at horizon we got stuck again (HBS), got that up and running and presets failing for MCSPU related reasons, working on it...

20:48 finally got over our issues (fixed a bad install of MCSPU software relative to TCS - LBT.conf wasn't updated on jet), and preset to BS9193 GS=0 to set IE=+22 CA=-63 , send active preset collimating /tmp wfsc 165-

Checking the Z22 coefficient

21:12 once collimated checking Z22 coefficient for Z4 compensation.

21:18 send active preset to BS9107 GS 0 little glitch, started guiding then suddenly stopped for no apparent reason (21:24), collimating wfsc 181-... so noticed large (500") Rx on m2 while doing focus 0 ran into M1 -y limit. Speculating that this is a transient thermal gradient effect, so watch for hitting -ve y limit on M1 when in positive thermal gradient.

21:38 reset IE=+15 CA=-61 , resend active preset. Notice as the guide image goes from a nice evenly illuminated ring wfsc 186 as the focus is driven out, some coma is clearly being added in (21:39). Refer to previous discussion about focus only correction for large initial aberration.

21:50 have large Z22 obvious in pupil, with +ve thermal gradient Z22 sign is reversed from sign last night in -ve thermal gradient. Quick adjustment of Z22 (+150 nm) with FCin set to zero. Great! No large focus driven in, will get to quantitative test once converged.

21:55 sent another +300 Z22 and see a bit of a jump in focus, maybe 500 nm.

21:57 Z22 now mainly gone 10-20 nm residuals, 450 nm sent.

22:00 now going to measure how much residual focus there is with Z22 correction, switching to IDL collimation..."STOP" WFS in GCS GUI. (Seeing ~ 2").

22:02 idl collimating wfsc 2-16 (note collimating with 15s integrations and converging rapidly despite low light and stray light from the nearby moon, switch to 25 s integrations for less jump on Z4). wfsc 9 shows evidence of vibration or pointing oscillation.

22:14 sending +200 nm Z22 and measuring Z4 wfsc unfortunately get bad new influence forces so clear C00 drive back in the 450 nm Z22 that I had before and recollimate wfsc 17- This time will drive -200 nm Z22 from here and measure focus...

22:21 do focus zero secondary odffload and pull back 1 mm from previous del Y adjustment. now 2 mm from -ve y limit, were 1 mm from it after last focus 0.

22:29 mirror panics gonig to Zenith to fix.

22:47 John kindly brings bucket near; collimated again, this time with only 300 nm Z22 loaded on as the higher number may have contributed to mirror panic

22:52 Now collimated sending -200 nm Z22 and measuring focus /AGW_Data wfsc 36-38 Z4 -581, -525, -475. Averaging and multiplying by scale factor 1.5 gives Z4 delta ~-800 nm when we send -200 nm Z22. Therefore deduce that 53 microns del Z per 100 nm of Z22 is too much by 15, testing hypothesis now, collimating wfsc-42-45

23:01 sent -200 nm got the delta Z adjustment and backed out 30 microns measuring wfsc 45-47, no measurable focus change, so the new and final amount of delta Z to go with 100 nm of Z22 is +37 microns.

John at helm for tracking tests

23:15 pointto BS9107 TRACK but don't see star on IRTC at J

23:18 pointto BS9107 ACQUIRE - Adjust IE=+15 CA=-66

23:20 pointto BS9108 ACQUIRE

23:26 accidentally sent the same ACQUIRE preset again, but this time it hung and didn't display the guide camera image.

23:28 stop and restart GCSL, but it refused to stop

23:29 pointto BS9108 TRACK

23:33 pointto BS9105 TRACK

23:35 pointto BS9105 gs=0 ACTIVE

23:38 focus 0 to offload corrections from M2

Switch to H filter

23:41 pointto BS9105 TRACK

23:46 takeseq 3.0 Nimage=500 take1+ with manual collimation update of M1 every 50 frames.

My eyeball impression is that we still have jumps at the 0.3 arcsec level. However, the seeing is not very good, and the focus drifted off in the latter part of the sequence, so this isn't a very impressive data set.

offset 30 30 to blank sky

00:13 takeseq 3.0 Nimage=50 take1+ for blank sky

But IRTC was set to write to R:/IRTC_Data/081113 while the directory was named 081113w, so maybe we didn't get this data?

00:17 pointto BS9105 gs=3 ACTIVE

Over to Doug for more tracking data

12:43 Continue collimating through gcs 250

12:45:45 takeseq 1 600 (10 minutes of tracking data with guiding and collimation) 57 elevation

12:44 collimation applied at the following IRTC frame: 39, 89, 138, 186, 237, 291, 347, 403, 462, 512, 556

12:55:30 done through gcs 266 File is /Repository/IRTC_Data/active_00075_cube.fit

12:55 Jesper joins us

Jesper data

1:05 IDL preset to M38coll star in ACQUIRE, Wayne set IE=+20 CA=-73

1:05 seeing 1.2"-1.5" from wfsc

1:14 preset to M38coll in ACTIVE

1:15 GCS couldn't adjust WFSC time below 15 sec. did not apply any correction. stop WFSing

1:19 start IDL agw_collimate_loop, EXP=10 (saturated SH spots) wfsc 49-58

1:25 preset to M38 in TRACK

1:31 setxy -x 0 -y 612.5, setfocus -p 31.6

1:32 readGuideCam 100 (500 ms) Can see all 5 stars. Offset 0 -8 to put constellation on center of chip

1:35 setExpTime 10000 to get good image ~141-168 (then got dim and bad seeing)

2:00 preset to M34 in ACTIVE gcs (dim on wfsc ~400 counts) 286-294

2:04 preset to M34 in TRACK mode

2:05 readGuideCam 100 (10 sec exp) guider 177-236

2:15 preset to BS9107 in ACTIVE (Wayne had to unwrap the cable wrap ~few minutes)

2:30 collimate gcs 295-303

2:39 preset BS9107 ACTIVE

39 226.92 0, 476.7 76
69 226.92 -87.584, 550.815 gstar not found

2:48 preset to M5_0162 TRACK to tweak IE=10 and CA=-75

2:52 preset to BS9107 ACTIVE Failed! could see bright star on acquisition image, but no guiding

2:53 preset BS9107 ACQUIRE tweak to IE=+20 CA=-70

2:57 preset to bs9107 ACTIVE Success gcs 304-306

gstar PA x,y fits cube exp frames
8 274.59 0.176, 432.321 jesper_00077_cube.fit 1.0 60
47 274.59   jesper_000_cube.fit 1.0 60

Note: bright star on acquire000100, 102, 103, 104 but did not find and guide.

3:10 preset BS9107 ACTIVE Succss collimate 310-314

gstar PA x,y fits cube exp frames
8 274.59 0.124, 432.351 jesper_00078_cube.fit (bad) 1.0 60

3:22 ~6940 Huge amount of focus and astigmatism: Several actuators on the primary go "wonky". Wayne reset them, but GCS has put ~2000 nm of correction on primary. Must re-collimate

2:25 preset BS9107 collimate gcs 318-325

gstar PA x,y fits cube exp frames
47 274.59 111.59, 551.055 jesper_00080_cube.fit 1.0 30
20 274.59 56.112, 451.65 jesper_00081_cube.fit 1.0 30
8 274.59 0.166, 432.416 jesper_00082_cube.fit 1.0 30
39 274.59 -100.106, 521.13 jesper_00083_cube.fit 1.0 30

4:04 preset BS9107 ACTIVE gstar=0 gcs 326-329 (330 and 331 on gstar=47)

gstar PA x,y fits cube exp frames
47 335.73 0.337, 485.016 jesper_00084_cube.fit 1.0 30
82 335.73 100.09, 534.312 jesper_00085_cube.fit 1.0 30
90 335.73 137.67, 581.253 jesper_00086_cube.fit 1.0 30
20 335.73 -113.502, 485.592 jesper_00087_cube.fit 1.0 30

4:36 IDL preset to M67 collimate gcs 332-336 guider 280 -307

4:48 IDL preset to M67 collimate 337-339

4:51 Stopguiding, move guide probe out by hand. See one star on IRTC. offset -9 8 and see three stars: jesper_00088_cube.fit and jesper_00089_cube.fit (background)

4:59 IDL preset to M67 ACTIVE gcs 340-341

5:02 preset M67 ACTIVE gcs 342

5:14 preset M57 ACTIVE PA=30 gcs 343-344

5:16 stop Guiding, move out probe, take 1secX60 IRTC: jesper_00090_cube.fit

5:22 takeseq 3secx10 jesper_00091_cube.fit (background jesper_00092_cube.fit)

5:30 My laptop freezes because hard drive is full. Transfer some data to lbtdu16

5:45 preset M67 ACTIVE PA=60 gcs 345-348

5:48 stop guiding move probe out, could not see any stars on IRTC

6:01 Done for tonight.

-- JohnHill - 12 Nov 2008
Topic revision: r8 - 14 Nov 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback